# Watermark Anything with Localized Messages

### v1

@@ Conclusion.
In spite of the JND weighting, we notice that the watermark can still be visible in some areas of the watermarked images, even at a relatively high PSNR (see examples in App. 12). This could be due to the JND map focusing only on the cover image and not on the watermark signal itself. Improvements might be achieved by employing more sophisticated Human Visual System (HVS) models or by regularizing the watermark to eliminate repetitive patterns during training.
@@
Section 5 provides the specifics of our training and evaluation setup. Further implementation details, including network architectures, mask designs, transformation parameters for robustness evaluation, and settings for baseline comparisons, are presented in App. 10. The models and code for inference and training used in this paper will be made available upon publication. The code is based on PyTorch, the training dataset is publicly available, and the training requires half a week on 8 V100 GPUs, which makes the models reproducible at reasonable cost.

### v2

@@ Conclusion.
In spite of the JND weighting, we notice that the watermark can still be visible in some areas of the watermarked images, even at a relatively high PSNR (see examples in App. F). This could be due to the JND map focusing only on the cover image and not on the watermark signal itself. Improvements might be achieved by employing more sophisticated Human Visual System (HVS) models or by regularizing the watermark to eliminate repetitive patterns during training.
Post-ICLR rebuttal changes.
We added Appendix G to further discuss WAM’s operating point and include additional experiments: more baseline comparisons, robustness to diffusion-based purification, neural compression, overlapping watermarks, and balancing the perceptibility/robustness trade-off.
@@
Section 5 provides the specifics of our training and evaluation setup. Further implementation details, including network architectures, mask designs, transformation parameters for robustness evaluation, and settings for baseline comparisons, are presented in App. D. The models and code for inference and training used in this paper will be made available upon publication. The code is based on PyTorch, the training dataset is publicly available, and the training requires half a week on 8 V100 GPUs, which makes the models reproducible at reasonable cost.